Day Shift summary - 6/28

ramberg (ramberg@fnal.gov)
Mon, 28 Jun 1999 15:32:21 -0500 (CDT)

Shift Summary for hhh shift, dd/mmm/99

On Shift: Val Prasad, Erik Ramberg

Hours of beam delivered at all intensities: 4
Hours of beam delivered at nominal intensity: 4
Hours of E832 data (special runs not included): 4
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam problems as MCR tunes to NUMI target
and attempts to raise intensity.

========================================================================
Runs ended this shift:
Run 13750 - 23.6M events - BEAM_832
Run 13751 - 7.4M events - BEAM_832

Tapes brought to FCC this shift: QK*052-54,QKB018

========================================================================
Accesses during shift: none

========================================================================
Description of Shift:

07:00 Take over from owls.
07:20 Val arrives. Wrists are slapped for tardiness
07:30 Run 13750 stops because tapes are full.
07:45 Start run 13751 with new tape (QK*055-57,QKB019)
08:00 Rick verifies that problem in EWUD logic is due to
bad discriminator in trigger system. (fixed this shift)
08:30 I notice that the 'sdaq_online' script has an error of
2 in the rates and inform Theo of this. (fixed this shift)
09:30 I deliver tapes to FCC. Due to inexplicable burst of kindness,
I also make a run to the highrise to pick up pancakes for Val.
11:00 Plane 2, cpu 6 has abnormally long processing time.
11:10 MCR calls to say there are problems with beam and we will get
reduced intensity for a few spills.
11:15 Beam position is off and MCR contacted.
11:30 MCR says they will play with tune. We stop run 13751. Last
20 spills are bad.
12:00 Rick replaces bad 4413 discriminator. Start run 13753 without
beam.
13:00 wait
14:00 wait
15:00 wait
15:20 MCR opens beam stop and states intention to raise slowly to 8E12.
For run 13753, spills 1-->142 are bad.
15:30 Hand off.